OpenStudy (anonymous):

Simplify. Assume r is greater than zero. √110 / 18r^10

OpenStudy (luffingsails):

Hi @talita , the easiest way to approach this is by consider only the numbers first. Trying simplifying those and see if you can make sense of it.

OpenStudy (luffingsails):

\[\sqrt{\frac{ 110 }{ 18r^{10} }} = \sqrt{\frac{ 55 }{ 9r^{10} }}\]

OpenStudy (anonymous):

thank you